John's age is 4 more than 3 times Michael's age. If the sum of their ages is 20, how old are John and Michael?
Answer by fractalier(6550)

Call them j and m.
j+m=20 and j=4+3m
Plug this second one into the first and get
4+3m + m = 20
4+4m=20
4m=16
m=4
and
j = 16